### v1

@@ 5.1 Model training methodology
While fine-tuning based on CoT has been used to augment datasets for solving scientific problems [25], its application to visualization tasks remains unexplored. Building upon the methodology by Ko et al. [19], we have devised a custom CoT strategy for the V-RECS problem domain, involving three ’Teachers,’ each tasked with addressing specific challenges related to the V-RECS tasks, including visualization recommendation and captioning, explanation of the process of column selection in T and filtering, and suggestion of additional questions for data exploration. The Teachers’ tasks are delineated in Table 1, with variations based on the input and intended goal.
@@
Finally, the responses generated by the three Teachers, along with the original query Q and the expected visualization V are combined according to a custom prompting strategy used to fine-tune the Student, as detailed next.
Table 1: Description of the Teacher’s tasks
@@
Explain to the user why the features used in
VegaZero are the most appropriate given all
@@
T2 VegaZero Describe Describe the content of the visualization
@@
suggest to the user for exploring the data
5.2 Implementation details
@@
In this section we provide the relevant implementation details to replicate our methodology. Additional details can be found in the annexed Supplementary Material and at the V-RECS GitHub  [^10] https://github.com/lucapodo/V-RECS.git.
@@
Visualization representation A challenging aspect concerns the choice of a visualization description method. A visualization can be represented as a declarative specification, a bitmap image, or through its code. The first type is based on visualization grammars, like the popular Vega-Lite. The second type considers the visualization as an image, and the third type represents the visualization as a set of instructions, for example, using Matplotlib in Python. Pixel-based visualizations would add limitations and complexity to our approach, as LLMs are primarily trained using textual data, while limiting interactivity and deployability. Similarly, employing a code-based visualization method would hinder the creation of a pure LLM approach, requiring a specific code environment to render and generate visualizations. Differently, a grammar-based visualization can be considered agnostic to the used rendering environment, making it more general than code, while still formatted as structured text. For this reason, we chose grammar-based visualizations as the most suitable for our scenario. Specifically, we have opted for VegaZero, proposed in [28].
@@
Selection of the training dataset Various datasets have been proposed in the literature for Natural Language to Visualization (NL2VIS) tasks, with NvBench [27] emerging as the most prominent and widely adopted. However, it exhibits limitations such as a lack of complexity in the visualization types, overly explicit natural language queries (except for some instances pre-classified as ”hard"), and the absence of support for visualization narratives or reasoning steps. While other datasets like NLVCorpus [44] or Quda [12] provide alternatives, they also come with their own constraints. In particular, NLVCorpus includes only 893 visualization-query pairs, and Quda does not provide the chart associated with the queries. In the absence of a dataset capable of supporting all our requirements, we opted to choose NVBench as the best fit for our purpose. We chose the NvBench version provided by NcNet, as it aligns closely with our task requirements, providing the queries, visualization, and data sources. Specifically, this version excludes visualizations generated from multiple data tables and adds VegaZero representations tailored to each query. Other datasets, such as VisText [48] or VIS30k[3], lack the query-vis-data triples and focus solely on caption generation.
@@
Teacher and Student prompting GPT-4 has been selected as the Teacher and prompted with three distinct tasks (T1, T2, and T3), described at a high-level in Table 1, based on the input and task delivered. Furthermore, leveraging the work of Ko et al. [19], we provide each prompt with in-context reference to the VegaZero template to help the model generate consistent visualization. Each task is tailored for Zero-shot CoT supplemented with strategically embedded sub-tasks, denoted here as ”steps". These steps act as ”guiding principles", offering the Teacher subtle cues on approaching and unraveling the task. The steps vary for each teacher, based on their task.
Figure 2 shows the detailed description of the prompt template for T1 (Explain), split into three parts and related steps. The other prompt templates, along with a more detailed description, can be inspected in the Supplementary Material.
@@
Figure 2: Structure of the T1 prompt. The related task is described in Table 1. It has three parts: A encourages the LLM to reason ”step-by-step" as per the Chain of Thoughts method, B includes the description of sub-tasks (steps) to guide the model, and C is the response template.
Next, all the teachers’ responses $R_{i}$ are combined into a prompt template to instruct the Student, consisting of the header, the input, the body, and the response, as follows:
@@
    The header includes the explicit task the model is supposed to pursue. It drives the model to understand what it has to achieve. Along with the task, this part also includes the VegaZero template structure and, finally, the "Let’s think step by step" formula [20] which is the basis of the Chain of Thought prompting method.
@@
    The body consists of the intermediate or reasoning steps the model should follow to select proper visualization properties, such as the aggregation function. This part is crucial to provide additional information to the end-user so that they can understand the model’s design choices and generate the final response. Adding this part enables the explainability of the generated response.
@@
    The response is the last part of the prompt, where we provide the structure of the answer that must be presented by the model to the end-user. It includes the VegaZero specification V to generate a chart, and the E,C,S triple that makes up the narrative explaining the recommended visualization.
The Student prompt template, along with a filled example of training instance, is depicted in Figure 3. The left-hand side of Figure 3 shows how the different components of the Teachers’ responses contribute to filling the student prompt.
Figure 3: Fine-tuning prompt template (Student prompt). The leftmost box shows how the responses R1, R2, and R3 contribute to filling the different parts of the template. The rightmost box is a (partly) filled template for a specific training/testing instance.
Fine-tuning The last step is fine-tuning the model. As a student, we have used Llama-2-7B [51], the smallest from the Llama2 family. To fine-tune Llama-2-7B, we generated a dataset comprising 2938 triples, each consisting of a dataset D, a user’s query Q, and a VegaZero specification V, following the same distribution of query complexity of the original dataset. Drawing from the QLora technique [9], our training approach went beyond attention modules [52] to incorporate linear components. QLora has shown very promising results during training. With reference to [9], when configuring Lora, we opted for an $r$ value of 64 and an $\alpha$ value of 128. During fine-tuning, we employed a batch size of 4, a learning rate of $1e^{-4}$, and utilized the AdamW optimizer.
@@
Figure 4 shows an example of V-RECS response at inference time, when given a dataset D and query Q.
6 Evaluation
@@
This section presents a comprehensive evaluation of V-RECS, encompassing quantitative and qualitative assessments. In both types of evaluation, we compare the results generated by V-RECS with the ones generated by GPT-4. While a baseline evaluation was initially planned with Llama-2-7B (the un-tuned model), extensive experimentation with prompt templates and incremental prompt construction yielded unsatisfactory results, leading to its exclusion from further evaluation.
Figure 4: Example of V-RECS response given a dataset and a user’s query
The evaluation was carried out using EvaLLM [35]. Similar to the popular ISO-OSI stack, EvaLLM introduces a layered evaluation framework where each layer’s output serves as the input for the layer above it, including five layers. The lower layers focus on the structural aspects of the visualization, while the upper layers delve into more semantic and insightful properties.
@@
[…]
    The code layer serves as the foundational level, evaluating fundamental structural properties.
@@
[…]
    The representation layer addresses core visualization properties based on established representation rules.
@@
[…]
    The presentation layer assesses the perceptual quality of data presentation.
@@
[…]
    The informativeness layer gauges the intrinsic quality of visualizations in terms of insights and adherence to best practices in visualization literacy.
@@
[…]
    Lastly, the LLM layer evaluates the computational and strategic costs of generating specific visualizations.
Furthermore, as detailed by Podo et al. [35], the EvaLLM stack introduces twelve different levels within each layer, each evaluating distinct properties.
Since EvaLLM is a framework, it must be specialized for specific experiments. For the purpose of evaluating V-RECS we chose to focus only on seven of the twelve levels proposed in EvaLLM. Additionally, we introduced a new level, ”Caption quality". We implemented the seven selected levels as follows:
[…]
Syntax correctness: For this level, we employed the Altair Python library’s render to validate VegaLite specifications generated by GPT and a parser developed by VegaZero authors for V-RECS. The evaluation metric used to compare correct specifications against the total evaluated examples was accuracy.
[…]
Note that all the above levels can be evaluated quantitatively, except for Significance, which we assessed with a qualitative evaluation by human judges.
6.1 Quantitative evaluation
For the quantitative evaluation, we used 1800 samples drawn from the original ncNet dataset [28], not used in the training phase, augmented with the generated visualization narratives, as described in Section 5. This experiment aimed to prompt both GPT-4 and V-RECS models with user queries and datasets, subsequently evaluating the responses generated by each model. These responses encompassed the recommended visualizations V and the accompanying narrative, allowing for a comprehensive performance assessment. In detail, the experimental setup involved prompting the models with a user query, dataset, and task description. Conversely, the prompt for GPT-4 was more elaborate, including the task, zero-shot CoT formulation, a set of tasks guiding its response generation, and the desired response format121212i.e., we use a template format similar to those used for the three teachers, see Section 5.2. As shown in Figure 5, the two models exhibited comparable performance across these levels, with GPT-4 showcasing a slight advantage from +3% to +4% accuracy in what is comparable performances between the two models. However, a notable discrepancy emerged in assessing axes quality, where GPT-4 appeared to significantly underperform compared to V-RECS (75% against 93% accuracy). Specifically, GPT-4 frequently exhibited errors such as axis inversions, a phenomenon previously discussed in literature [35]. Despite this, both models demonstrated comparable accuracy across the first six levels, indicating comparable efficacy in generating meaningful visualizations and narratives. Furthermore, it is essential to consider the efficiency aspect of our evaluation. Efficiency, depicted for all dimensions in Figure 5 (right), entails comparing the accuracy achieved by each model at each level, while also considering the model complexity. Notably, our findings suggest that results comparable to GPT-4 can be achieved with a much smaller and more manageable model, which is a relevant advantage in practical deployment scenarios.
Figure 5: Results of the EvaLLM evaluation experiment
6.2 Qualitative evaluation
For the Significance layer, we have designed a qualitative experiment involving 50 randomly selected samples (not used for training) evaluated by 15 human annotators. Importantly, we maintained the original query difficulty distribution (as categorized in the source dataset), ensuring that our evaluation was representative of real-world scenarios. To facilitate the evaluation process, we developed a web application designed specifically for conducting a comparative study  [^13] https://vrecseval.streamlit.app/. This application allowed us to present each volunteer with pairs of responses from the two models for the same query and dataset. The identity of the models’ responses was hidden from participants, ensuring unbiased evaluations. Each participant evaluated the responses using a Likert scale, ranging from 1 (Completely Meaningless) to 5 (Completely Meaningful). This scale provided a structured framework for capturing the nuanced differences in the quality and relevance of the responses. Firstly, each participant was assigned a subset of 20 samples from the total pool of 50. This ensured that each sample received multiple evaluations from different participants, thereby mitigating individual biases. To minimize the impact of any outliers or inconsistencies in the evaluations, we enforced a minimum threshold of three unique evaluations per sample.
Figure 5, in the Significance layer, shows a slight advantage of V-RECS. This is better shown in Figure 6, where results are distributed according to the complexity of the query, and categorized into four levels as in the original dataset. As expected, the quality of answers worsens with the complexity of natural language queries for both systems, with V-RECS performing slightly better except for the extra hard sentences.
Figure 6: Results of the qualitative evaluation by complexity of the natural language query
@@ 6.2 Qualitative evaluation
Upon a subsequent models’ evaluation by the authors, we noticed that the explanations and captions produced by the two systems are slightly different but equally satisfactory, although V-RECS provides a more comprehensive explanation of the reasoning steps. This result mitigates some of the concerns reported in the user study by Kim et al. [18] in terms of lack of depth and critical thinking capabilities while explaining a visualization design. Contrarily, GPT-4 often surpasses V-RECS in suggesting additional questions, presenting greater diversity and serendipity. We believe this difference can be reduced in our future research by improving the fine-tuning prompt strategy for the T3 task. We also noted a couple of hallucinations in GPT-4, while in two cases V-RECS was able to produce a good narrative but, due to some problem in handling data values, could not generate a visualization.
7 Discussion and Future work
This paper has proposed the first LLM-based visual recommender, V-RECS, capable of providing explanations, captioning, and suggestions for further analyses, and has demonstrated its capabilities. Remarkably, V-RECS is based on fine-tuning a ”small", open-source model, Llama-2-7B, while achieving comparable performance to a much larger, proprietary model, GPT-4.

### v2

@@ V-A Model training methodology
While fine-tuning based on CoT has been used to augment datasets for solving scientific problems [42], its application to visualization tasks remains unexplored. Building upon the methodology by Ko et al. [40], we have devised a custom CoT strategy for the V-RECS problem domain, involving three “Teachers,” each tasked with addressing specific challenges related to the V-RECS tasks, including visualization recommendation and captioning, explanation of the process of column selection and filtering in D, and suggestion of additional questions for data exploration. The Teachers’ tasks are delineated in Table II, with variations based on the input and intended goal.
@@
Finally, the responses generated by the three Teachers, along with the original query Q and the expected visualization V are combined to fine-tune the Student, as detailed below.
TABLE II: Description of the Teacher’s tasks
@@
Explain to the user why the features used in the
VegaZero specification V are the most appropriate, given all
@@
T2 VegaZero Caption Generate a caption to describe the content of the visualization
@@
suggest to the user for further exploring the data
V-B Implementation details
@@
In this section we provide the relevant implementation details to replicate our methodology. Additional details can be found in the annexed Supplemental Material and in the V-RECS GitHub  [^8] https://github.com/lucapodo/V-RECS.git.
@@
Visualization representation A challenging aspect concerns the choice of a visualization description method. A visualization can be represented as a declarative specification, a bitmap image, or through its code. The first type is based on visualization grammars, like the popular Vega-Lite. The second type considers the visualization as an image, and the third type represents the visualization as a set of instructions, for example, using Matplotlib in Python. Pixel-based visualizations would add limitations and complexity to our approach, as LLMs are primarily trained using textual data, while limiting interactivity and deployability. Similarly, employing a code-based visualization method (such as some of the systems surveyed in Table I) would hinder the creation of a pure LLM approach, requiring a specific code environment to render and generate visualizations. Differently, a grammar-based visualization can be considered agnostic to the used rendering environment, making it more general than code, while still formatted as structured text. For this reason, we chose grammar-based visualizations as the most suitable for our scenario. Specifically, we have opted for VegaZero, proposed in [12].
Selection of the training dataset Various datasets have been proposed in the literature for Natural Language to Visualization (NL2VIS) tasks, with NvBench [43] emerging as the most prominent and widely adopted. The natural language queries in the NvBench dataset were generated by starting with existing NL2SQL benchmarks, modifying SQL queries into visualization queries, filtering out low-quality visualizations, and then manually verifying the quality of the synthesized (NL, VIS) pairs. Despite its widespread adoption, the dataset exhibits limitations such as a lack of complexity in the visualization types, overly explicit natural language queries (except for some instances pre-classified as “hard”), and the absence of support for visualization narratives or reasoning steps. Additionally, due to the generation process described above, some of the NL queries in the dataset may not appear natural999as shown by some of the examples used throughout this paper, which are all taken from the NvBench dataset. However, this apparent problem replicates a well-studied [44] real-world challenge, where occasional and/or non-expert users often employ informal, ungrammatical, and sketchy language. Recall that one of the motivations for leveraging natural language understanding features in visual recommenders is to target these types of users. Thanks to their advanced language processing capabilities, LLM4VIS may infer a user’s information need even when expressed in complex, incomplete, or sketchy form.
@@
While other datasets like NLVCorpus [45] or Quda [46] provide alternatives, they also come with their own constraints. In particular, NLVCorpus includes only 893 visualization-query pairs, and Quda does not provide the chart associated with the queries. In the absence of a dataset capable of supporting our training requirements (i.e., sufficiently large samples of D,Q,V triples), we opted to choose NVBench as the best fit for our purpose. We chose the NvBench version provided by NcNet, as it aligns closely with our training task requirements, i.e., providing the queries, visualizations, and data sources. Specifically, this version excludes visualizations generated from multiple data tables and adds VegaZero representations tailored to each query. Other datasets, such as VisText [47] or VIS30k[48], do not include data-query-visualization (D,Q,V) triples and focus solely on caption generation.
@@
Teacher and Student prompting GPT-4 has been selected as the Teacher and prompted with three distinct tasks (T1, T2, and T3), described at a high-level in Table II, based on the input and task delivered. Furthermore, leveraging the work of Ko et al. [40], we provide each prompt with in-context reference to the VegaZero template to help the model generate consistent visualization. Each task is tailored for Zero-shot CoT supplemented with embedded sub-tasks, denoted here as “steps”. These steps act as “guiding principles”, offering the Teacher subtle cues on approaching and unraveling the task. The steps vary for each task. Figure 3 shows the detailed description of the prompt template for T1 (Explain), composed of three blocks and related steps. Note that the explanation is split into two parts, described by Steps 2 and 3 of Figure 3 (box B). The other prompt templates, along with a more detailed description, can be inspected in the Supplemental Material.
@@
Figure 3: Structure of the T1 prompt. The related task is described in Table II. It has three parts: A encourages the LLM to reason “step-by-step” as per the Chain of Thoughts method, B includes the description of sub-tasks (steps) to guide the model, and C is the response template. Note, in box B, that Step 2 and Step 3 direct the system to split the explanation into two parts. Part 1 (E1) summarizes the user’s information needs, to make sure they have been correctly interpreted; part 2 (E2) explains why, given D and Q, certain features X and Y and transformation functions have been selected.
Figure 4: Training template during fine-tuning. The leftmost box shows how the Teacher’s responses R1, R2, and R3 contribute to filling the different parts of a Student’s training instance. The rightmost box is a (partly) filled template for a specific training instance.
Figure 5: Example of V-RECS response at inference time for the query Which product lines generate the most revenue?
To instruct the Teacher, we used a dataset comprising 2938 NvBench triples, each consisting of a dataset D, a user’s query Q, and a VegaZero specification V, following the same distribution of query complexity of the original dataset101010the complexity level is provided by the creators of the dataset. . Based on the three prompts, the Teacher generates, for each input triple (D, Q, V) of the dataset, the related responses $R_{1}$, $R_{2}$, and $R_{3}$. Next, the three responses $R_{i}$ along with the corresponding D, Q, and V are combined into a training instance template used to instruct the Student, consisting of the header, the input, the body, and the response, as follows (see Figure 4):
@@
    The header includes the explicit task the model is supposed to pursue. It drives the model to understand what it has to achieve. Along with the task, this part also includes the VegaZero template structure.
@@
    The body consists of the intermediate reasoning steps the model should follow to select proper visualization properties, such as the aggregation function. This part is crucial to provide additional information to the end-users so they can understand the model’s design choices and generate the final response. Adding this part enables the explainability of the generated response.
@@
    The response is the last part of the prompt, where we provide the structure of the answer that the model must present to the end-user. It includes the VegaZero specification V to generate a chart (the recommended visualization), and the E,C,S triple that makes up the visualization narrative.
The training instance template, along with a (partially) filled training example, is depicted in Figure 4. The left-hand side of Figure 4 shows how the different components of the Teachers’ responses contribute to filling the template  [^11] More details on how the Teacher’s responses are combined to fill the template are provided in the Supplemental Material.. Further note in Figure 4, left, the partly ungrammatical structure of the query. The query shown is included in the NvBench dataset, and, as previously remarked, mimics the tendency of non-expert users to employ a sketchy language.
Fine-tuning As a result of the Teacher CoT phase, the initial 2938 triples are enriched with the visualization narratives generated by GPT, and packed into training instances, shaped as in Figure 4. The last step is fine-tuning the Student with the generated training instances.
As a Student, we have used Llama-2-7B [14], the smallest from the Llama2 family. Drawing from the QLora technique [49], our training approach went beyond attention modules [50] to incorporate linear components. QLora has shown very promising results during training. With reference to [49], when configuring Lora, we opted for an $r$ value of 64 and an $\alpha$ value of 128. During fine-tuning, we employed a batch size of 4, a learning rate of $1e^{-4}$, and utilized the AdamW optimizer.
@@
Figure 5 shows an example of V-RECS response at inference time. As shown, during inference, V-RECS is given a dataset D and query Q and returns a visualization recommendation along with a narrative composed of the three elements E, C, and S.
VI Evaluation
@@
This section presents a comprehensive evaluation of V-RECS, encompassing quantitative and qualitative assessments on evaluating V,E,C and S. The evaluation aims to ascertain to what extent V-RECS has achieved the proposed objectives, i.e. generating insightful visualization recommendations supported by rich narratives, while relying on a non-proprietary and low-cost LLM. In both quantitative and qualitative evaluation, we compare the results generated by V-RECS with the ones generated by GPT-4. While a baseline evaluation was initially planned with Llama-2-7B (the un-tuned model), extensive experimentation with prompt templates and incremental prompt construction yielded unsatisfactory results, leading to its exclusion from further evaluation. We emphasize that the only possible comparison is with GPT-4, the most powerful LLM to date since other LLM4Vis (and NL2Vis) systems do not produce visualization narratives, nor can they be easily adapted to do so. The ability to generate rich visualization narratives represents the main contribution of V-RECS compared to the state of the art, which would make a comparison with models without this functionality inappropriate.
[…]
To evaluate V-RECS, we focused only on four of the five layers originally proposed in EvaLLM, developing our stack implementation. Additionally, for the informativeness layer, we introduced a new level, “Narratives quality”. We implemented the selected levels as described in Table III. Note that the first four levels in Table III can be evaluated quantitatively, while both Informativeness and Effort are assessed with a qualitative evaluation.
TABLE III: Description of the implemented evaluation levels based on the EvaLLM framework. The layers to which every level belongs are indicated by the colored boxes. ”Ground-truth” refers to a subset of 1800 (D,Q,V) triples, generated as discussed in Section V, not used to train V-RECS.
Levels Metric Implementation
[…]
correctness
Grammar
accuracy
It employs the Altair Python library to validate GPT responses and a parser developed by VegaZero authors for V-RECS.
[…]
mapping
Data mapping
accuracy
It measures the model’s ability to select appropriate columns from dataset D based on user queries Q. The metric is computed by comparing the columns selected by each model against the ground-truth columns.
[…]
correctness
Mark
accuracy
This level computes the metric by comparing the mark selected by each model against the ground-truth.
[…]
quality
Axes
accuracy
This level was evaluated based on accuracy, comparing the predicted axis mapping against the ground-truth.
[…]
quality
Human
evaluation
This level involves a human-based qualitative evaluation in the form of a comparative study based on a custom web app (see Section VI-A).
[…]
Computational
cost
Evaluates the complexity of the strategy used to obtain the desired results, ranging from zero-shot prompting for GPT (low cost) to fine-tuning for V-RECS (high cost).
Figure 6: Results of the EvaLLM evaluation experiment
VI-A Quantitative evaluation
The quantitative evaluation involves Syntax correctness, Data mapping, Mark correctness, and Axes quality (see Table III). This experiment aimed to prompt both GPT-4 and V-RECS models with a user’s query and datasets, subsequently evaluating the responses generated by each model. These responses encompassed the recommended visualizations V and the accompanying narrative (E,C,S), allowing for a comprehensive performance assessment. For V-RECS the prompt solely consisted of the user query, dataset, and task description. Conversely, the prompt for GPT-4 was more elaborate, including the task, zero-shot CoT formulation, a set of sub-tasks guiding its response generation, and the desired response format121212i.e., we use a template format similar to those used for the three teachers, see Section V-B.
For the evaluation, we used 1800 samples drawn from the original ncNet dataset [12], not used in the training phase, augmented with the generated visualization narratives, as described in Section V. As shown in Figure 6, the two models exhibited comparable accuracy across three quantitative levels (Syntax, Data, and Mark), with GPT-4 showcasing a slight advantage from +3% to +4% accuracy. However, a notable discrepancy emerged in assessing Axes quality, where GPT-4 appeared to significantly underperform compared to V-RECS (75% against 93% accuracy). Specifically, GPT-4 frequently exhibited errors such as axis inversions, a phenomenon previously discussed in literature [20]. Despite this, both models demonstrated comparable accuracy, indicating comparable efficacy in generating meaningful visualizations.
VI-B Qualitative evaluation
Figure 7: Average rank of the qualitative evaluation on a 1-5 scale. The first histogram to the left evaluates the visualization V, the second shows the overall usefulness of the visualization narrative VN. The three subsequent histograms separately evaluate E, C, and S. For each type of narrative, participants have been asked to consider both the informativeness (i_score) and usefulness (u_score).
Figure 8: Narratives evaluation for the 16 non-expert participants. Results are confirmed even for this cohort, with a widening in performance in favor of V-RECS for the explanation narrative.
Qualitative evaluation involves the Informativeness and Effort layers. For the Informativeness layer, we have designed a qualitative experiment involving 50 randomly selected samples (not used for training) evaluated by 35 human annotators (27 males, 8 females, average age$=29.6$ years, $std=9.8$). It followed an initial study with 15 participants where we mainly assessed the quality of the generated visualizations and the supporting narratives altogether. We maintained the original query difficulty distribution (as categorized in the source dataset), ensuring that our evaluation was representative of real-world scenarios. In addition to 50 examples of (D,Q) pairs extracted from NvBench, we also added ten slightly more natural sounding and complex sentences, to test the system ”into the wild”. To facilitate the evaluation process, we developed a web application designed specifically for conducting a comparative study  [^13] https://vrecseval.streamlit.app/. This application allowed us to present each volunteer with pairs of responses from the two models for the same query and dataset. The identity of the models’ responses was hidden from participants, ensuring unbiased evaluations. Each participant was tasked to evaluate the generated visualizations to assess their quality; for each of the associated narratives (E, C, and S), the participant was asked to express its informativeness and utility in supporting the interpretation of the visualization. We finally asked for an overall judgment on the utility of the narratives as a whole. All the responses used a Likert scale, ranging from 1 (Completely Meaningless) to 5 (Completely Meaningful). This scale provided a structured framework for capturing the nuanced differences in the quality and relevance of the responses. Firstly, each participant was assigned a subset of 10 samples from the total pool of 60. This ensured that each sample received multiple evaluations from different participants, thereby mitigating individual biases. To minimize the impact of any outliers or inconsistencies in the evaluations, we enforced a minimum threshold of three unique evaluations per sample.
Figure 7 shows the results of the evaluation. Looking at the first two charts on the left, we note how GPT4 and V-RECS present comparable performance for both visualization generation quality and overall narrative quality. Interestingly, V-RECS is slightly outperformed by GPT-4 for visualization quality while slightly outperforming GPT-4 in narrative quality. The narratives analysis was further investigated individually, with results visible in the right part of figure 7. Interestingly, V-RECS outperforms GPT-4 for explanation quality in both informativeness (VRECS: $\mu=3.54$, $std=1.40$; GPT-4: $\mu=2.91$, $std=1.31$) and utility (VRECS: $\mu=3.19$, $std=1.49$; GPT-4: $\mu=2.74$, $std=1.38$). This tells us that V-RECS is more capable of explaining the generation process to a user, helping in fostering more trust in the resulting visualization. An inverted situation happens for the captioning narratives, with GPT-4 proving more capable than V-RECS at summarizing the content of a generated visualization. We believe that the generally richer captions generated by GPT-4 felt more informative and useful to the users. Due to a caption being a summary description of what a visualization shows, this result may be partially explained by the better capabilities of larger models like GPT-4 to produce effective summaries.
Finally, looking at the suggestions for further analysis narrative, V-RECS presents a slight edge on both perceived informativeness (VRECS: $\mu=3.37$, $std=1.18$; GPT-4: $\mu=3.16$, $std=1.23$) and utility (VRECS: $\mu=3.15$, $std=1.25$; GPT-4: $\mu=2.89$, $std=1.29$). Overall, the behavior on the S narrative presents comparable results between the two models. We further analyzed the narratives by looking at the sub-cohort of nonexpert users (15 out of 35 declaring a degree of expertise ranging from none:1 to average:3, where average means the knowledge of basic visualization techniques and their usage in current workflow analyses). Figure 8 shows the results. Performance of V-RECS and GPT-4 remains confirmed for narratives C and S, with just a slight increase for both models on all the scores. More interestingly, the differences in the Explanation quality are increased for both informativeness (VRECS: $\mu=3.78$, $std=1.09$; GPT-4: $\mu=2.98$, $std=1.11$) and utility (VRECS: $\mu=3.32$, $std=1.23$; GPT-4: $\mu=2.69$, $std=1.14$), still in favor of V-RECS. This result supports the capability of V-RECS to provide help to non-expert users in visualization for data analysis tasks. Finally, considering the potential effect of the query difficulty, the results of this study confirmed that the quality of answers worsens with the complexity of natural language queries for both models, with V-RECS performing slightly better except for the extra hard sentences where results are comparable between the models. For the sake of space, the evaluation by complexity levels is reported in the supplemental material, along with other data.
The overall results of the manual evaluation, reported in aggregated form, have been added to the Narrative importance level of the informativess layer, as depicted in Figure 6.
To conclude the qualitative analysis, we consider the LLM effort and the efficiency aspects of our evaluation. Efficiency, depicted for all dimensions in Figure 6 (right), entails comparing the accuracy achieved by each model at each level, while also considering the model complexity. Notably, our findings suggest that results comparable to GPT-4 can be achieved with a much smaller and more manageable model, which is a relevant advantage in practical deployment scenarios.
On the other hand, V-RECS’ implementation effort (see the LLM effort level in Figure 6) demands significantly more work compared to GPT. While GPT relies primarily on prompting techniques, V-RECS requires extensive fine-tuning, increasing the effort needed to achieve results comparable to GPT.
VI-C Error Analysis
Based on human evaluation and subsequent analysis by the authors, we detected common issues in V-RECS and GPT-4, comparing their outputs. Visualization errors in both systems can be categorized as follows:
@@ VI-C Error Analysis
    Incorrect Scaling: Both systems sometimes struggle to scale the axes correctly, resulting in unusual and confusing data presentations. For example, in Figure 9b, GPT-4 erroneously encoded years as numeric quantities, resulting in a misleading x-axis.
@@
    Inverted Axes: Sometimes, the models map columns to the incorrect axis, resulting in inverted axes. For example, V-RECS in Figure 9a shows such an error.
@@
    Non-Optimal Zoom/Spacing: Both systems fail to optimize the visualization zoom, leading to charts where data is overlapped in a small section while a large portion remains empty and white.
@@
    Hallucinations: GPT-4 sometimes generates misleading visualizations. As shown in Figure 9a, GPT-4 created a new bar chart version with no meaningful visual representation, violating standard visualization principles.
@@
    Missing Data: Both models occasionally generate correct Vega-Lite specifications but fail to map the data, resulting in an empty grid and visualization.
  6. 6.
    Input errors For a limited number of (Q,D) pairs, we noted that either Q is hardly interpretable even by a human, or missing/wrong data formats are present in D. This is due to problems in the NvBench dataset, already highlighted in Section V-B. Surprisingly, although these problems lead to wrong visualizations, both systems are often still able to interpret the user’s information need with substantially correct narratives.
All discussed errors are common to both models, except hallucination, which we detected only in GPT-4. We postulate that V-RECS is less prone to hallucinations due to its more straightforward grammar structure for generating the visualizations. VegaLite allows users to customize visualizations extensively, which can sometimes result in misleading visual representations. Conversely, V-RECS is built upon VegaZero, a streamlined version of VegaLite. VegaZero reduced the complexity by reducing the customization options, ensuring that the focus remains on the core properties of the visualization, thereby mitigating the risk of generating erroneous visual content.
The above-listed issues highlight existing challenges in ensuring the reliability of automated visualization systems and underscore the need for further improvements in their design and implementation.
Figure 9: Examples of errors in both systems. (a) Example of hallucination (GPT hallucinates). (b) Example of incorrect scaling (GPT) and spacing(V-RECS)
TABLE IV: Comparison of GPT-4 and V-RECS visualization narratives (VN) for the query:”What is the average rank of pilots in different positions?”. V-RECS narratives are colored in violet, and GPT-4 narratives are colored in blue.
VN Responses
E The ’position’ and ’rank’ columns are selected because they directly answer the user’s question. The ’position’ column provides the different categories (pilots) and the ’rank’ column provides the numeric data to be averaged. The ’mean’ aggregate function is used to calculate the average rank for each position. The ’group’ transform is applied to group the data by position, which allows us to calculate the average rank for each position. Other columns like pilot id, pilot name, age, nationality, join year, and team are not selected as they do not contribute to answering the user’s question. The ’rank’ and ’position’ features are the best among all the others from the dataset based on the user instruction because the user wants to know the average rank of pilots in different positions The ’rank’ feature provides the rank of the pilots and the ’position’ feature provides the different positions of the pilots
C The visualization is a bar chart where the x-axis represents the ’Position’ and the y-axis represents the ’Rank’. The height of each bar corresponds to the mean rank for each position. The bars are grouped by ’Position’. The visualization is a bar chart that shows the average rank of pilots in different positions. The x-axis represents the different positions and the y-axis represents the average rank. Each bar corresponds to a position and the height of the bar represents the average rank of pilots in that position
S 1) Which position has the highest average rank? 2)Which position has the lowest average rank? 3) Are there any positions where the average rank is significantly different from others? 1) What is the average age of pilots in different positions? 2)What is the average join year of pilots in different positions? 3)What is the distribution of ranks in different positions?
@@
As far as the E,C,S narratives are concerned, we noticed that the explanations and captions produced by the two systems are slightly different but equally satisfactory, with V-RECS producing richer explanations and GPT better captions, as also highlighted by the human evaluation in Section VI-B. Specifically, V-RECS provides a more comprehensive explanation of the reasoning steps as shown in the example of Table IV. This result mitigates some of the concerns reported in the user study by Kim et al. [51] in terms of lack of depth and critical thinking capabilities while explaining a visualization design. On the other hand, despite judges have considered V-RECS suggestions slightly more useful to explore the recommended visualization, we noted that GPT-4 suggestions go beyond the user’s query, presenting greater diversity and serendipity, as shown in the example of Table IV. We believe this feature (adding suggestions for additional exploration of the dataset rather than solely on the presented visualization) could be implemented in our future research by improving the fine-tuning prompt strategy for the T3 task.
VII Conclusions and Future work
This paper has proposed V-RECS, the first LLM-based visual recommender enhanced with narratives, and has demonstrated its capabilities. The proposed model mitigates, among others, two relevant challenges of state-of-the-art LLM4Vis systems, discussed in Section I: lack of verifiability and controllability. Concerning the first, V-RECS allows both expert and non-expert users to evaluate the quality and significance of a proposed visualization through a rich narrative composed of an explanation and a caption. In addition, it also suggests additional queries for further data exploration. Concerning the second challenge, V-RECS is based on fine-tuning a “small”, open-source model, Llama-2-7B, while achieving comparable performance to a much larger, proprietary model, GPT-4, as we demonstrated in Section VI.